Compute the annual rate of inflation in the GDP deflator between the two periods (the circled data below the highlighted quarter and year). Please show your work.


Table 1.1.6. Real Gross Domestic Product, Chained Dollars Billions of chained (2012) dollarsl Seasonally adjusted at annual rates ast Revised on: December 21, 2018 -Next Release Date January 30, 2019 2017 2017 2017 2017 2018 201 Q4 18,223.8 18,324.0 18,5 Q1 17,863.0 17,995.2 1 12,427.6 12,5159 12,584.9 12,706.4 12,722.8 12,8 4,307.3 4,366.0 4,410.2 4,483.9 4,477.0 4, 1,527.2 1,559.2 1,588.6 1,636.6 1,628.2 1 2,786.3 2,813.9 2,829.9 2,857.7 2,858.6 2, Gross domestie product 8,120.8 2 Personal consumption expenditures 3 Goods Durable goods Nondurable goods 6 Services
1.6. Real Gross Domestic Product, Chained Dollars llonesised on: December 21, 2018- Next Release Date January 30, 2019 ined (2012) dollars] Seasonally adjusted at annual rates 2018 Q3 18,665.0 12,953.3 4,585.5 1,677.4 2,919.2 Line Gross domestie product 2 Personal consumption expenditures 3 Goods Durable goods 5 Nondurable goods
Answer #1

Assume 2017 as base year

for year 2017 =( 18120.8/18120.8 )*100= 100

for 2018 = (18665/18120.8)*100 = 103

Annual inflation rate = (value 2018/value 2017)-1)*100

Annual inflation rate = ((103/100)-1)*100

Annual inflation rate = (1.03-1)*100 = 3%